OpenStudy (anonymous):

A vending machine offers 8 different drinks. One day, 6 employees each purchased a drink from the vending machine. Find the probability that at least 2 employees purchased the same drink. Round your answer to the nearest hundredth.

OpenStudy (qqstory):

this is hard

OpenStudy (anonymous):

I'm bad with stuff like this.

OpenStudy (qqstory):

1-6/8

OpenStudy (qqstory):

do you have the answer?

OpenStudy (anonymous):

total cases: each man has 8 choices.so total cases = 8^6. each has different drink: cases = 8*7*6*5*4*3. for at least two have same drink = total cases - when each have different drink = 8^6 - 8*7*6*5*4*3. probability = (8^6 - 8*7*6*5*4*3) / (8^6).

OpenStudy (anonymous):

Okay, I got it: First, you have a probability of getting 1 type of drink on a vending machine with 8 drinks: \[P (Drink 1) = (1\div8)\] but since the even requires at least 2 of the 6 employees to choose the same drink, we have: \[P(Drink 1) = (1\div8) \times (6 employees \div 2 chosesamedrink)\] \[P(1) = (1\div8)\times(6\div2)\] Hope that helps, all you have to do now is to work from the right, divide 6 to two, and multiply it by 1/8. (Then simplify).

OpenStudy (anonymous):

ok i finally got it thanks !!!

OpenStudy (anonymous):

The first guy drinks a drink. It can be anything. The second guy has a (7/8) chance of drinking something else. Now 2 different drinks have been drank. The third guy has a (6/8) chance of drinking something else. Fourth guy: (5/8) fifth guy: (4/8) sixth: (3/8) Probability they all drink something different: 1*(7/8)*(6/8)*(5/8)*(4/8)*(3/8)

OpenStudy (anonymous):

The probability that at least two drink the same thing will be 1-(that probability^)

OpenStudy (anonymous):

smooth got it, go with that also remember that "at least two" or "at least one" is a set up for "not one" or "not two"

OpenStudy (anonymous):

But realistically, that's not a very good model of the probability. It assumes equally likely choices, and coke is way more popular than anything else.

OpenStudy (anonymous):

coke adds life